# ProSGNeRF: Progressive Dynamic Neural Scene Graph with Frequency Modulated Auto-Encoder in Urban Scenes

### v2

@@ 4 Experiments
In this section, we validate the proposed ProSGNeRF method on the existing automotive datasets KITTI [8] and VKITTI [3]. We validate that our method outperforms existing dynamic representation methods in image reconstruction, novel view synthesis, and dynamic object manipulation.
Datasets and Metrics. We evaluate ProSGNeRF on Kitti tracking datasets [8] for the real-world urban scene. The Kitti dataset [8] is captured using a stereo camera. Each sequence consists of more than 100 time steps and images of size $1242\times 375$, each from two camera perspectives and more than 12 dynamic objects from different classes. We also use synthetic data from the Virtual KITTI 2 Dataset [3]. We report the standard view synthesis metrics PSNR, SSIM, and LPIPS scores for all evaluations.
Implementation details and Baselines. ProSGNeRF is evaluated on a desktop PC with NVIDIA RTX 3090ti GPU. For experimental settings, the auto-encoder uses the ResNet34 backbone pre-trained on ImageNet. Each decoder uses the five fully-connected residual blocks. We set $\lambda_{c}=1$ , $\lambda_{d}=0.005$ ,$\lambda_{\sigma}=0.005$, $\lambda_{seg}=0.001$. The bounding box dimensions are scaled to include the shadows for better reconstruction. For further details of our implementation, refer to the supplementary. We compare our method against state-of-art implicit neural rendering methods: the original NeRF implementation [14], NeRF+Time, NSG [16], and SUDS [21].
@@ 4.1 Experimental Results
Image Reconstruction and Novel View Synthesis. We conduct image reconstruction and novel view synthesis experiments in urban dynamic scenes. In Tab. 1, we present our qualitative results of image reconstruction and view synthesis on KITTI datasets [8]. We evaluate the methods using different train/test splits, holding out the full dataset, every fourth time step, every other time step, and finally, training with only one in every four time steps. We present quantitative results with the PSNR, SSIM, and LPIPS metrics. The proposed method outperforms all baseline methods on all metrics. Especially as the number of training views decreases, our method demonstrates a more pronounced advantage in novel view synthesis.
@@
The qualitative results are shown in Fig. 3. This figure depicts the appearance of the dynamic scene at different time instances from left to right (time step: 66, 80, 102, 234). The vanilla NeRF implicit rendering fails to reconstruct dynamic scenes accurately. It can only represent static scenes, which leads to nearly identical render frames for all the sequences. The NSG algorithm provides a relatively accurate representation of dynamic objects but cannot handle scenes with camera ego-motion. This limitation arises from the fixed root of their scene graph representation, which cannot move with the camera. In the last column of the figure, their reconstruction results appear highly blurred when the camera moves. SUDS exhibits poor representation capability for dynamic objects; their objects appear blurry in Fig. 1. Especially as the number of training views decreases, all the existing methods generate ghosting artifacts near areas of movement. We also present object PSNR and LPIPS curves as the training view falls in Fig. 5. ProSGNeRF achieves the best reconstruction and novel view synthesis results in dynamic object representation and large-scale camera ego-motion.
Methods VKITTI [3]-50%
NeRF NeRF + Time NSG SUDS Ours
[…]
Table 3: Novel View Synthesis on VKITTI datasets [3]. We outperform past works on image reconstruction accuracy, following their experimental protocol and self-reported accuracy.
@@
Scene Decomposition and Object Manipulation. Compared to other methods, our approach excels significantly in its ability for scene decomposition and object manipulation. The structure of the proposed method can naturally decompose scenes into dynamic and static scene components. We present the rendering of isolated nodes of our progressive scene graph in Fig. 4 to present the scene editing ability. In Fig. 4, we compared the scene decomposition ability of our method and other methods. We can see that our method learns a clean decomposition between the objects and the background.
@@
We then generate previously unseen frames of novel object arrangements, temporal variations. Our progressive neural scene graph architecture allows dynamic urban scene editing, which can be manipulated at its edges and nodes. To exhibit the flexibility of our proposed scene representation, we manipulate the pose of dynamic objects through rotation. We present the view synthesis results in Fig. 4. Note that we have not observed the rotation of dynamic objects during training.
@@
In addition to object manipulation and node removal of the dynamic objects, our method allows for constructing completely novel scene graphs and novel view synthesis. Fig. 4 shows view synthesis results that we randomly insert new object nodes and new object positions with new transformations. We constrain samples to the joint set of all observed road trajectories. Note that all these translations and manipulations have not been observed during training.
@@ 4.2 Ablation Study
We conduct an ablation study to verify the importance of the components of our method. The results of our experiments are presented in Tab. 4.
@@
Frequency auto-encoder. In the second experiment, we replace our frequency auto-encoder with the original MLP. We can see that the network cannot effectively represent dynamic objects due to the sparsity and movement of dynamic objects.
@@
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o progressive scene graph 27.21 0.782 0.079
w/o frequency auto-encoder 28.32 0.823 0.069
w/o depth loss 28.93 0.824 0.078
w/o KL divergence loss 29.45 0.84 0.073
Full Model 30.31 0.891 0.055
Table 4: Ablation study. We conduct experiments on KITTI tracking 0006 sequence [8] to verify the effectiveness of our method. Our full model achieves the best view synthesis results.
@@ 5 Conclusion
Limitations  Although we present a first attempt at building city-scale dynamic environments, many problems still need to be solved. We need the ground truth camera pose, which is challenging to get in the real-world environment.
6 Acknowledgement
Authors gratefully appreciate the contribution of Yanbo Wang from Shanghai Jiao Tong University, Zirui Wu and Yongliang Shi from Institute for AI Industry Research (AIR), Tsinghua University.

### v3

@@ 5 Experiments
In this section, we validate the proposed ProSGNeRF method on the existing autonomous driving datasets KITTI [kitti], VKITTI [vkitti], Waymo [waymo] and nuScenes datasets [nuScenes]. We validate that our method outperforms existing dynamic representation methods in image reconstruction, novel view synthesis, and dynamic object manipulation.
Figure 11: We illustrate the scene editing results for the single object movement setting. As shown in the figure, our method enables flexible insertion of different types of vehicles into the scene, including trucks and cars.
@@ 5.1 Experimental Results
Image Reconstruction and Novel View Synthesis. We conduct image reconstruction and novel view synthesis experiments in urban dynamic scenes. Table 3 reports the qualitative results of image reconstruction and novel view synthesis on the KITTI dataset [kitti]. To comprehensively evaluate performance under varying data availability, we adopt multiple train and test splits, including using the full dataset, holding out every fourth time step, holding out every other time step, and training with only one out of every four time steps. Quantitative evaluations are performed using PSNR, SSIM, and LPIPS metrics. Across all experimental settings, the proposed method consistently outperforms all baseline approaches on all evaluation metrics. Compared with NeRF based methods, our approach achieves substantially improved performance in both background reconstruction and dynamic object reconstruction. This improvement can be attributed to the progressive scene graph design, which enables large-scale ego-motion scene modeling, as well as the incorporation of multiresolution hash encoding in the background network, allowing more accurate representation of fine scale geometric and texture details. In comparison with 3DGS based methods, our approach demonstrates superior performance in dynamic object reconstruction and exhibits stronger generalization capability. Specifically, our method is able to effectively model a wide range of dynamic object categories across different scenes, highlighting its robustness and adaptability in complex dynamic environments. In particular, as the number of available training views decreases, our method exhibits a more significant advantage in novel view synthesis performance.
@@
In addition to KITTI, we also present image reconstruction and novel view synthesis results on the VKITTI [vkitti], Waymo [waymo], and nuScenes [nuScenes] datasets in Tab. 3, Tab. 4, Tab. 5, and Tab. 6. The experimental results across these diverse urban driving benchmarks clearly demonstrate the effectiveness and strong generalization capability of the proposed method, showing its ability to accurately model both static background structures and dynamic objects in complex real world scenes. The qualitative results are shown in Fig. 5 and 7. Figure 5 depicts the appearance of the dynamic scene at different time instances from left to right (time step: 66, 80, 102, 234). The vanilla NeRF implicit rendering fails to reconstruct dynamic scenes accurately. It can only represent static scenes, which leads to nearly identically rendered frames for all the sequences. The NSG algorithm provides a relatively accurate representation of dynamic objects but cannot handle scenes with large-scale camera ego-motion. This limitation arises from the fixed root of their scene graph representation, which cannot move with the camera. In the last column of the figure, their reconstruction results appear highly blurred when the camera moves. SUDS exhibits poor representation capability for dynamic objects; their objects appear blurry in Fig. 5. Methods based on 3DGS achieve relatively high accuracy in background modeling. However, they exhibit limited performance in dynamic object reconstruction, suffer from poor generalization capability, and rely heavily on high precision point cloud initialization such as LiDAR measurements. In contrast, our method leverages prior knowledge encoded in a foundation model to guide scene understanding, enabling effective extraction of object appearance and shape features. This design allows the proposed approach to robustly handle the sparse observation problem caused by high speed object motion, while simultaneously maintaining strong generalization performance across different scenes and object categories. Especially as the number of training views decreases, all the existing methods generate ghosting artifacts near areas of movement. We also present object PSNR and LPIPS curves as the training view falls in Fig. 6. ProSGNeRF achieves state of the art performance in dynamic object reconstruction and novel view synthesis under large scale camera ego motion, which strongly validates the effectiveness of the proposed frequency modulated foundation model design.
Table 7: Ablation study of our design choices on the full 0006 sequence of KITTI datasets [kitti] for image reconstruction and 50% for novel view synthesis. The metrics are PSNR, SSIM, LPIPS.
Methods Image Reconstruction Novel View Synthesis-50%
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o progressive scene graph 26.34 0.849 0.079 23.98 0.795 0.137
w/o overlapping between local scene graph 28.79 0.862 0.065 24.98 0.834 0.103
with only MLP 28.75 0.865 0.065 24.81 0.832 0.104
with only Dino v2 29.41 0.881 0.060 25.98 0.865 0.088
with frequency modulation and MLP 27.18 0.860 0.068 24.72 0.831 0.105
w/o object pose input 27.98 0.864 0.071 24.75 0.822 0.103
with shared encoders for shape and appearance 27.15 0.867 0.073 24.28 0.816 0.113
using the shared network for car and truck 27.81 0.877 0.071 24.56 0.839 0.101
w/o depth loss 29.25 0.885 0.063 25.19 0.865 0.094
w/o KL divergence loss 29.28 0.891 0.069 25.45 0.860 0.096
w/o far-field loss 29.37 0.890 0.067 25.28 0.862 0.095
without planes sampling 29.68 0.896 0.065 25.89 0.864 0.091
without ray-box sampling 29.72 0.897 0.063 25.82 0.865 0.090
Full Model 30.31 0.931 0.057 26.55 0.878 0.086
@@
Scene Decomposition and Object Manipulation. Compared to other methods, our approach excels significantly in its ability for scene decomposition and object manipulation. The structure of the proposed method can naturally decompose scenes into dynamic and static scene components. We present the rendering of isolated nodes of our progressive scene graph in Fig. 8 and Fig to present the scene editing ability. In Fig. 8, we compared the scene decomposition ability of our method and other methods. We can see that our method learns a clean decomposition between the objects and the background.
@@
We then generate previously unseen frames of novel object arrangements, temporal variations. Our progressive neural scene graph architecture allows dynamic urban scene editing, which can be manipulated at its edges and nodes. To demonstrate the flexibility of the proposed scene representation, we rotate the pose of dynamic objects. We present the view synthesis results in Fig. 8 and Fig. 11. Note that we have not observed the rotation of dynamic objects during training.
@@
In addition to object manipulation and node removal of the dynamic objects, our method allows for constructing completely novel scene graphs and novel view synthesis. Fig. 8 shows view synthesis results that we randomly insert new object nodes and new object positions with new transformations. We constrain samples to the joint set of all observed road trajectories. Note that all these translations and manipulations have not been observed during training. Figure 11 illustrates the scene editing results for the single object movement setting. As shown in the figure, our method enables flexible insertion of different types of vehicles into the scene, including trucks and cars. These results demonstrate the strong object level editability of the proposed approach and highlight the advantages of the scene graph structure in supporting controllable and semantically consistent scene manipulation.
@@ 5.2 Ablation Study
In this section, we provide more detailed ablation studies of our method. The detailed experimental results is shown in Tab. 7. The table presents an ablation study showing the impact of different design choices on the performance of a scene representation model. It evaluates image reconstruction and novel view synthesis on the KITTI dataset across three metrics: PSNR, SSIM, and LPIPS.
Several variants of the model are compared by incrementally removing or changing components. This includes ablations of the progressive scene graph, loss functions, encoder sharing schemes, input modalities, and sampling strategies. The bottom row shows results for the full model with all components enabled.
The full model with all features performs the best across most evaluation metrics. The most impactful design choices are the progressive scene graph, separate encoders for shape and appearance, and ray-based sampling. Removing or changing these degrades performance, demonstrating their importance. Other ablations produce smaller differences, suggesting less critical roles. Moreover, we present the ablation study on progressive scene representation with pre-dividing in Tab. 8.
Overall, the ablation quantitatively validates design decisions through comparative analysis. It highlights key model architectures and training techniques that boost reconstruction fidelity and view synthesis photorealism. The table effectively summarizes component importance across multiple metrics.
Table 8: Ablation study on KITTI [kitti] and Waymo [waymo] datasets. We demonstrate the effectiveness of progressive scene representation and online scaling.
Methods KITTI [kitti] Waymo [waymo]
PSNR SSIM LPIPS PSNR SSIM LPIPS
pre-divide 24.89 0.827 0.118 31.28 0.908 0.112
w/o scaling 25.29 0.837 0.108 32.85 0.915 0.103
Ours 25.95 0.856 0.095 33.80 0.931 0.092
@@
Frequency modulated foundation model. In the second experiment, we replace our frequency foundation model with the original MLP. We can see that the network cannot effectively represent dynamic objects due to the sparsity and movement of dynamic objects.
@@ 6 Conclusion
Availability of data and materials All the datasets used in the paper will be publicly available.